# On Minimal Achievable Quotas in Multiwinner Voting

### v1

@@ B.3 Restricted Domains
Step. Let the claim hold for $j-1\geq 0$. We show that it also holds for $j$. If $c_{j}$ is not selected by the algorithm, then $C_{j}=C_{j-1}\subseteq W^{(j-1)}$ by the induction hypothesis. Therefore, assume our algorithm selects $c_{j}$. Since $c_{j}$ is selected by the algorithm, there is an $\alpha$-JR violation when considering $C_{j-1}\cup\{c_{j+1},\dots,c_{m}\}$ by a group of voters $N^{\prime}$ approving a common candidate $c$ with $c\sqsubseteq c_{j}$ and no voter $v\in N^{\prime}$ approves a candidate $c^{\prime}\in\{c_{j+1},\dots,c_{m}\}$. By assumption, $W^{(j-1)}$ does not have any violations, which means there is some candidate $c^{\prime\prime}\in W^{(j-1)}$ which is approved by at least one voter $v\in N^{\prime}$ and $c^{\prime\prime}\sqsubseteq c_{j}$. If $c^{\prime\prime}=c$, the statement trivially holds, so assume in the following that $c\neq c^{\prime\prime}$. Now, we denote as $W^{(j)}$ the committee after replacing $c^{\prime\prime}$ in $W^{(j-1)}$ with candidate $c_{j}$ and claim that $W^{(j)}$ does not violate $\alpha$-JR. For this, assume for the sake of contradiction that there exists a cohesive group $(N^{\prime},c)$ that causes an $\alpha$-JR violation for $W^{(j)}$. This implies that no voter in $N^{\prime}$ approves a candidate in $W^{(j)}$. First, consider the case $c_{j}\sqsubset c$. Since $W^{(j-1)}$ satisfies $\alpha$-JR there exists a $v\in N^{\prime}$ with $c^{*}\in A_{v}\cap W^{(j-1)}$ for some $c^{*}$. $c^{*}\in W^{(j)}$ would imply that $i$ approves a candidate in $W^{(j)}$. Thus, $c^{*}=c^{\prime\prime}$. But since $v$ approves $c$ and $c^{*}$ and $c^{*}\sqsubset c_{j}\sqsubset c$ this implies that $i$ approves $c_{j}$; a contradiction.
@@
Thus, in the following, we can assume that $c\sqsubset c_{j}$. By definition of the algorithm, we know that the subset of candidates $C_{j-1}\cup\{c_{j},\dots,c_{m}\}$ satisfies $\alpha$-JR. Therefore, there exists a voter $v\in N^{\prime}$ that approves a candidate $c^{*}\in C_{j-1}\cup\{c_{j},\dots,c_{m}\}$, but no candidate in $W^{(j)}$. Since $C_{j-1}\cup\{c_{j}\}\subseteq W^{(j)}$ it has to hold that $c_{j}\sqsubset c^{*}$. But then voter $v$ approves both $c^{*}$ and $c$, which together with $c\sqsubset c_{j}\sqsubset c^{*}$ implies that $c_{j}\in A_{v}$. Since we have a CI instance, this implies that $v$ approves $c_{j}$, which contradicts our assumption that $v\in N^{\prime}$. This concludes the proof. ∎
@@ C.1 Model Configuration
For the experiments, we use the Python voting library by Lackner et al. (2023). We generate approval ballots under two commonly used preference models: Impartial culture (IC) and the Euclidean Threshold model. For both models, we vary the number of voters $n\in\{11,12,29,59\}$, the number of candidates $m\in\{5,9,15\}$, and the committee size $k\in\{3,5,8,11\}$. Under IC, each voter independently approves candidates with probability $p\in\{0.3,0.5\}$. Under the Euclidean model, candidates and voters are embedded in the plane with voters sampled around the origin $(0,0)$ using a Gaussian distribution with $\sigma=0.5$, and approvals are determined by a distance threshold $t\in\{1.7,2.3\}$ (see (Lackner et al., 2023) for more details). For each of the considered voting rules, we compute at most five committees. For the computation of the performance of each voting rule, we compare the optimal $\alpha$-value with the mean $\alpha$-value of the returned committees.
@@ C.2 Further Results on $\alpha$-EJR
We complement the results on $\alpha$-EJR presented in Section˜7 with Table˜1. We separate the performance of the voting rules according to the parameters $k,m,$ and $n$. Every entry is averaged over 400 randomly generated instances. This comes from the fact that for every $(k,m,n)$-tuple, we have two sampling configurations for both sampling models.
PAV performs the best out of the considered methods. MES and sequential Phragmén still perform reasonably well. For all combinations of parameters, the average additive distance between the $\alpha$-value achieved by the voting rules and the optimal $\alpha$-value is between 0.01 and 0.2, showing that these three rules are actually not as bad as the worst-case would suggest. This is especially surprising for sequential Phragmén, which does not satisfy 1-EJR. Unsurprisingly, CC performs badly relative to the other rules as its only concern is to maximize the number of voters that approve a candidate. Nevertheless, CC often contains committees that also satisfy EJR, resulting in an overall moderate performance in approximating the optimal $\alpha$-value. (Bredereck et al., 2019). Therefore, we need to note that if one were not to consider the worst of the committees returned by CC, it would have performed better in our analysis.
@@ C.3 Plots for JR
We present the same plots that we have considered for EJR, for JR. The distribution of the $\alpha$-values is significantly closer towards 0 than for EJR. The fact that the optimal value is 0 in almost $80$ per cent of the instances in the IC model shows that there are many instances in which committees satisfying every voter with at least one candidate exist. Figure˜8 again shows the performance of the four voting rules we considered in this analysis. CC is very close to the optimal $\alpha$-value for both the IC and EUT model. This is not surprising, since CC maximizes the coverage, and therefore has an optimal $\alpha$-value of 0 whenever $\alpha_{\text{JR}}^{*}$ is 0.

### v2

@@ B.3 Restricted Domains
Step. Let the claim hold for $j-1\geq 0$. We show that it also holds for $j$. If $c_{j}$ is not selected by the algorithm, then $C_{j}=C_{j-1}\subseteq W^{(j-1)}$ by the induction hypothesis and we can conclude the claim by choosing $W^{(j)}=W^{(j-1)}$. Therefore, assume our algorithm selects $c_{j}$. Further, if $c_{j}\in W^{(j-1)}$ it holds that $C_{j}\subseteq W^{(j-1)}$ and again the claim follows. Since $c_{j}$ is selected by the algorithm, there is an $\alpha$-JR violation when considering $C_{j-1}\cup\{c_{j+1},\dots,c_{m}\}$ by a group of voters $N^{\prime}$ approving a common candidate $c$ with $c\sqsubseteq c_{j}$ and no voter $v\in N^{\prime}$ approves a candidate $c^{\prime}\in\{c_{j+1},\dots,c_{m}\}$. By assumption, $W^{(j-1)}$ does not have any violations, which means there is some candidate $c^{\prime\prime}\in W^{(j-1)}$ which is approved by at least one voter $v\in N^{\prime}$ and $c^{\prime\prime}\sqsubseteq c_{j}$. Since we assume that $c_{j}\not\in W^{(j-1)}$ we can conclude that $c_{j}\neq c^{\prime\prime}$. Now, we denote as $W^{(j)}$ the committee after replacing $c^{\prime\prime}$ in $W^{(j-1)}$ with candidate $c_{j}$ and claim that $W^{(j)}$ does not violate $\alpha$-JR. For this, assume for the sake of contradiction that there exists a cohesive group $(N^{\prime\prime},c)$ that causes an $\alpha$-JR violation for $W^{(j)}$. This implies that no voter in $N^{\prime\prime}$ approves a candidate in $W^{(j)}$. First, consider the case $c_{j}\sqsubset c$. Since $W^{(j-1)}$ satisfies $\alpha$-JR there exists a $v\in N^{\prime\prime}$ with $c^{*}\in A_{v}\cap W^{(j-1)}$ for some $c^{*}$. $c^{*}\in W^{(j)}$ would imply that $v$ approves a candidate in $W^{(j)}$. Thus, $c^{*}=c^{\prime\prime}$. But since $v$ approves $c$ and $c^{*}$ and $c^{*}\sqsubset c_{j}\sqsubset c$ this implies that $v$ approves $c_{j}$; a contradiction.
@@
Thus, in the following, we can assume that $c\sqsubset c_{j}$. By definition of the algorithm, we know that the subset of candidates $C_{j-1}\cup\{c_{j},\dots,c_{m}\}$ satisfies $\alpha$-JR. Therefore, there exists a voter $v\in N^{\prime\prime}$ that approves a candidate $c^{*}\in C_{j-1}\cup\{c_{j},\dots,c_{m}\}$, but no candidate in $W^{(j)}$. Since $C_{j-1}\cup\{c_{j}\}\subseteq W^{(j)}$ it has to hold that $c_{j}\sqsubset c^{*}$. But then voter $v$ approves both $c$ and $c^{*}$, which together with $c\sqsubset c_{j}\sqsubset c^{*}$ implies that $c_{j}\in A_{v}$. This contradicts our assumption that $v\in N^{\prime\prime}$ and thus concludes the proof.
The runtime complexity of the proposed algorithm lies in $O(m^{2}n\log n)$. We need to check at most $n$ distinct $\alpha$-values according to Lemma˜4.10 and the algorithm iterates over all $m$ alternatives, checking in each step whether there is an $\alpha$-JR violation, which can be done in $O(mn)$. Using binary search for the $\alpha$-values results in an additional factor of $\log n$. ∎
@@ C.1 Model Configuration
For the experiments, we use the Python voting library by Lackner et al. (2023). We generate approval ballots under two commonly used preference models: Impartial culture (IC) and the Euclidean Threshold model (EUT). For both models, we vary the number of voters $n\in\{11,12,29,59\}$, the number of candidates $m\in\{5,9,15\}$, and the committee size $k\in\{3,5,8,11\}$. Under IC, each voter independently approves candidates with probability $p\in\{0.3,0.5\}$. Under the Euclidean model, candidates and voters are embedded in the plane with voters sampled around the origin $(0,0)$ using a Gaussian distribution with $\sigma=0.5$, and approvals are determined by a distance threshold $t\in\{1.7,2.3\}$ (see (Lackner et al., 2023) for more details). For each voting rule, we compute at most 5 committees. To compute each voting rule’s performance, we compare the optimal $\alpha$-value with the mean $\alpha$-value of the returned committees.
@@ C.2 Further Results on $\alpha$-EJR
We complement the results on $\alpha$-EJR presented in Section˜7 with Table˜2. We separate the performance of the voting rules according to the parameters $k$, $m$, and $n$. Every entry is averaged over 400 randomly generated instances. This comes from the fact that for every $(k,m,n)$-tuple, we have two sampling configurations for both sampling models.
PAV performs best among the considered methods. MES and sequential Phragmén still perform reasonably well. For all combinations of parameters, the average additive distance between the $\alpha$-value achieved by the voting rules and the optimal $\alpha$-value is between 0.01 and 0.2, showing that these three rules are actually not as bad as the worst-case would suggest. This is especially surprising for sequential Phragmén, which does not satisfy 1-EJR. Unsurprisingly, CC performs poorly relative to the other rules, as its sole concern is maximizing the number of voters who approve a candidate. Nevertheless, CC often contains committees that also satisfy EJR, resulting in an overall moderate performance in approximating the optimal $\alpha$-value. (Bredereck et al., 2019). Therefore, we need to note that if one were not to consider the worst committees returned by CC, our analysis would have performed better.
@@ C.3 Plots for JR
We present the same plots we considered for EJR for JR. The distribution of the $\alpha$-values is significantly closer towards 0 than for EJR. The fact that the optimal value is 0 in almost $80$ per cent of the instances in the IC model shows that there are many instances in which committees exist that satisfy every voter with at least one candidate. Figure˜8 again shows the performance of the four voting rules we considered in this analysis. CC is very close to the optimal $\alpha$-value for both the IC and EUT model. This is not surprising, since CC maximizes the coverage, and therefore has an optimal $\alpha$-value of 0 whenever $\alpha_{\text{JR}}^{*}$ is 0.